# Backward Conformal Prediction

### v3

@@ B.2 Introductory setting: binary classification
The adaptive miscoverage $\tilde{\alpha}$ as described in Equation (5) is computed using binary search with tolerance 0.005. For each run, we record the corresponding value of $\tilde{\alpha}$ and the resulting prediction set $\hat{C}_{n}^{\tilde{\alpha}}(X_{\text{test}})$. In Figure 6, we plot the histogram of the $N$ values of $1-\tilde{\alpha}$ obtained across runs in green, with the mean $1-\mathbb{E}[\tilde{\alpha}$] shown as a green dashed line. We also report the empirical coverage probability $\mathbb{P}(Y_{\text{test}}\in\hat{C}_{n}^{\tilde{\alpha}}(X_{\text{test}}))$ as a red dashed line. We observe that the coverage probability $\mathbb{P}(Y_{\rm test}\in\hat{C}_{n}^{\tilde{\alpha}}(X_{\rm test}))$ consistently exceeds the marginal coverage $1-\mathbb{E}[\tilde{\alpha}$], providing additional empirical support for the Taylor approximation underlying the marginal guarantee (2).
@@ B.2.1 Adaptive size constraint rule
We observe that the coverage probability exceeds the empirical coverage, supporting the validity of the Taylor approximation used to derive (2). Additionally, as in the constant case, the estimator $\hat{\alpha}^{\text{LOO}}$ increasingly approximates the marginal miscoverage $\mathbb{E}[\tilde{\alpha}$] as $n$ grows, in line with Theorem 3.5. These observations suggest that Backward Conformal Prediction provides a principled approach to uncertainty quantification in machine learning yielding controlled-size prediction sets, even when the set size is dynamically adapted based on the data.
@@ C.2 General case
Appendix D Deriving Marginal Coverage from the Post-Hoc Guarantee without Approximation
In this section, we derive a guarantee conceptually aligned with the coverage guarantee in (2), but obtained more directly, without relying on a Taylor approximation, and fully estimable from the calibration data. We present the proof only for the constant-size constraint rule, as in Theorem 3.1, but the same reasoning naturally extends to the more general setting of Theorem 3.5.
Theorem D.1.
Fix $\delta>0$ and assume that $\mathbb{E}[\tilde{\alpha}^{2}]>\eta$ for some $\eta>0$. Let
$\tilde{R}_{\delta}(n):=S_{\max}\frac{\sqrt{\frac{\log(4/\delta)}{2n}}+\frac{2}{n}}{\mu\left(S_{\min}/S_{\max}-\frac{1}{n}\right)}+\sqrt{\frac{\log(4/\delta)}{2n}}+\frac{2S_{\max}^{2}}{\mu S_{\min}}\frac{n+1}{n}\sqrt{\frac{\pi}{2(n+1)}}$
be the upper bound obtained in Theorem 3.1. Let $n_{\delta}$ be such that for all $n\geq n_{\delta}$,
$2\tilde{R}_{\delta}(n)<\mathbb{E}[\tilde{\alpha}^{2}]-\eta,$
which is well defined since $\lim_{n\rightarrow\infty}\tilde{R}_{\delta}(n)=0$ and $\mathbb{E}[\tilde{\alpha}^{2}]-\eta>0$. Suppose further that the score function $S$ is bounded and takes values in $S_{\min},S_{\max}]$, with $0~<~S_{\min}~\leq~S_{\max}$, and that the number of calibration samples satisfies $n~>~\min(n_{\delta},S_{\max}/S_{\min},\frac{\log(4/\delta)}{2}(\frac{S_{\max}}{\mu})^{2})$. In addition, assume that the vectors $\mathbf{E^{\rm test}}$, $\mathbf{E}^{j}$, $\mathbf{\tilde{E}^{\rm test}}$, and $\mathbf{\tilde{E}}^{j}$ satisfy the following properties almost surely:
  1. (P1)
  2. (P2)
    For all $y\in\mathcal{Y}$, $\mathbf{E}_{y}\neq 1$.
Then, if the samples $(X_{i},Y_{i})$ are i.i.d., we have with probability at least $1-\delta$:
$\mathbb{P}(Y_{\rm test}\not\in\hat{C}_{n}^{\tilde{\alpha}}(X_{\rm test}))\leq\left(\!\sqrt{\hat{\alpha}^{\rm LOO}_{\rm sq}}\!+\!\tilde{R}_{\delta}(n)/\sqrt{\eta}\!\right)\!\left(\frac{n+1}{n}\right)\!\frac{\sqrt{\mathbb{E}[S(X,Y)^{2}]}}{\!\mathbb{E}[S(X,Y)]\!-\!S_{\max}\sqrt{\frac{\log(4/\delta)}{2n}}\!+\!\frac{1}{n}S_{\min}},$
where
$\hat{\alpha}^{\rm LOO}_{\rm sq}:=\frac{1}{n}\sum_{j=1}^{n}\tilde{\alpha}_{j}^{2}.$
This means that for sufficiently large calibration sets, the miscoverage probability is effectively controlled by $\sqrt{\hat{\alpha}^{\rm LOO}_{\rm sq}}\frac{\sqrt{\mathbb{E}[S(X,Y)^{2}]}}{\mathbb{E}[S(X,Y)]}$, with the remainder term decreasing at a rate of $1/\sqrt{n}$ as the calibration size increases.
@@ Appendix D Deriving Marginal Coverage from the Post-Hoc Guarantee without Approximation
We start by rewriting the miscoverage probability:
[…]
$=\mathbb{E}\left[\mathbb{1}_{\{\tilde{\alpha}E^{\rm test}\geq 1\}}\right]$
$\leq\mathbb{E}[\tilde{\alpha}E^{\rm test}]$
$\leq\sqrt{\mathbb{E}[\tilde{\alpha}^{2}]}\,\sqrt{\mathbb{E}[(E^{\rm test})^{2}]},$
where the last inequality follows from the Cauchy-Schwarz inequality. We now bound the two factors separately.
Concerning the first factor, we first state a corollary from Theorem C.3, which follows directly from the fact that $x\mapsto x^{2}$ is 2-Lipschitz on $[0,1]$:
Corollary D.2.
$\lvert F(\mathbf{E})^{2}-F(\mathbf{\tilde{E}})^{2}\rvert\leq 2\lVert\mathbf{E}-\mathbf{\tilde{E}}\rVert$,
for all $\mathbf{E},\mathbf{\tilde{E}}\in\mathcal{K}$.
Using Corollary D.2 and mimicking the proof of Theorem 3.1, we obtain
$\lvert\hat{\alpha}^{\rm LOO}_{\rm sq}-\mathbb{E}[\tilde{\alpha}^{2}]\rvert=\left\lvert\frac{1}{n}\sum_{j=1}^{n}F(\mathbf{E}^{j})^{2}-\mathbb{E}[F(\mathbf{E}^{\rm test})^{2}]\right\rvert\leq 2\tilde{R}_{\delta}(n).$ (12)
By the assumption $2\tilde{R}_{\delta}(n)<\mathbb{E}[\tilde{\alpha}^{2}]-\eta$, it follows that $\hat{\alpha}^{\rm LOO}_{\rm sq}>\eta$. Hence, both $\hat{\alpha}^{\rm LOO}_{\rm sq}$ and $\mathbb{E}[\tilde{\alpha}^{2}]$ lie in the interval $(\eta,1]$. Since $x\mapsto\sqrt{x}$ is Lipschitz on $(\eta,1]$ with constant $1/(2\sqrt{\eta})$, we obtain
$\left|\sqrt{\hat{\alpha}^{\rm LOO}_{\rm sq}}-\sqrt{\mathbb{E}[\tilde{\alpha}^{2}]}\right|\leq\frac{\tilde{R}_{\delta}(n)}{\sqrt{\eta}}.$
In particular, this gives
$\sqrt{\mathbb{E}[\tilde{\alpha}^{2}]}\leq\sqrt{\hat{\alpha}^{\rm LOO}_{\rm sq}}+\frac{\tilde{R}_{\delta}(n)}{\sqrt{\eta}}.$
We now bound the second factor $\sqrt{\mathbb{E}[(E^{\rm test})^{2}]}$. Applying the same concentration inequality that was used to bound $\frac{1}{n}\sum_{j=1}^{n}S(X_{j},Y_{j})$ when establishing (12), we obtain
$(E^{\rm test})^{2}\leq\frac{S(X_{\rm test},Y_{\rm test})^{2}}{\bigl(\frac{n}{n+1}\bigr)^{2}\left(\mu-S_{\max}\sqrt{\frac{\log(4/\delta)}{2n}}+\frac{1}{n}S_{\min}\right)^{2}}.$
Taking expectations yields
$\mathbb{E}[(E^{\rm test})^{2}]\leq\frac{\mathbb{E}[S(X_{\rm test},Y_{\rm test})^{2}]}{\bigl(\frac{n}{n+1}\bigr)^{2}\left(\mu-S_{\max}\sqrt{\frac{\log(4/\delta)}{2n}}+\frac{1}{n}S_{\min}\right)^{2}}.$
Combining the bounds for both factors yields the claimed probabilistic bound. ∎
The assumption $\mathbb{E}[\tilde{\alpha}^{2}]>\eta$ ensures that we can apply a Lipschitz bound for the square-root function on $(\eta,1]$, while the conditions on $n$ simply guarantee that all the quantities we manipulate are well-defined and that no divisions by zero occur.
Note that $\mathbb{E}[S(X,Y)]$ can be estimated by the empirical mean $\frac{1}{n}\sum_{i=1}^{n}S(X_{i},Y_{i})$, and similarly $\mathbb{E}[S(X,Y)^{2}]$ by $\frac{1}{n}\sum_{i=1}^{n}S(X_{i},Y_{i})^{2}$. Applying an additional union bound then yields a high-probability miscoverage bound that can be computed from the calibration set.

### v4

@@ B.2 Introductory setting: binary classification
The adaptive miscoverage $\tilde{\alpha}$ as described in Equation (4) is computed using binary search with tolerance 0.005. For each run, we record the corresponding value of $\tilde{\alpha}$ and the resulting prediction set $\hat{C}_{n}^{\tilde{\alpha}}(X_{\text{test}})$. In Figure 6, we plot the histogram of the $N$ values of $1-\tilde{\alpha}$ obtained across runs in green, with the mean $1-\mathbb{E}[\tilde{\alpha}$] shown as a green dashed line. We also report the empirical coverage probability $\mathbb{P}(Y_{\text{test}}\in\hat{C}_{n}^{\tilde{\alpha}}(X_{\text{test}}))$ as a red dashed line. We observe that the coverage probability $\mathbb{P}(Y_{\rm test}\in\hat{C}_{n}^{\tilde{\alpha}}(X_{\rm test}))$ consistently exceeds the marginal coverage $1-\mathbb{E}[\tilde{\alpha}$], providing additional empirical support for the coverage guarantee (2).
@@ B.2.1 Adaptive size constraint rule
We observe that the coverage probability exceeds the empirical coverage, supporting the validity of (2). Additionally, as in the constant case, the estimator $\hat{\alpha}^{\text{LOO}}$ increasingly approximates the marginal miscoverage $\mathbb{E}[\tilde{\alpha}$] as $n$ grows, in line with Theorem 3.5. These observations suggest that Backward Conformal Prediction provides a principled approach to uncertainty quantification in machine learning yielding controlled-size prediction sets, even when the set size is dynamically adapted based on the data.
@@ C.2 General case
Appendix D Proof of Proposition 2.2
For completeness, we give here a proof of Proposition 2.2, which can be found in [Koning, 2025b]:
Proposition D.1.
Consider a calibration set $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and a test data point $(X_{\rm test},Y_{\rm test})$ such that $(X_{1},Y_{1}),\dotsc,(X_{n},Y_{n}),(X_{\rm test},Y_{\rm test})$ are exchangeable. Let $\tilde{\alpha}>0$ be any miscoverage level that may depend on this data. Then, for the conformal set
$\hat{C}_{n}^{\tilde{\alpha}}(x):=\left\{y:\frac{S(x,y)}{\frac{1}{n+1}\left(\sum_{i=1}^{n}S(X_{i},Y_{i})+S(x,y)\right)}<1/\tilde{\alpha}\right\}$,
the coverage inequality (2) holds, i.e.,
$\mathbb{P}(Y_{\rm test}\in\hat{C}_{n}^{\tilde{\alpha}}(X_{\rm test}))\geq 1-\mathbb{E}[\tilde{\alpha}$].
@@
Since $E^{\rm test}$ is an e-value (that does not depend on $\alpha$), we have $\mathbb{E}[\sup_{\alpha>0}\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}/\alpha]\leq 1$, because $\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}=\mathbb{1}\{\alpha E^{\rm test}\geq 1\}\leq\alpha E^{\rm test}$. Rewriting this inequality gives $1/\mathbb{E}[\sup_{\alpha>0}\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}/\alpha]\geq 1$, and applying Jensen’s inequality then yields $\mathbb{E}[1/(\sup_{\alpha>0}\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}/\alpha)]\geq 1$, which is equivalent to $\mathbb{E}[\inf_{\alpha:E^{\rm test}\geq 1/\alpha}\alpha]\geq 1$, or, equivalently, $\mathbb{E}[1-\inf_{\alpha:E^{\rm test}\geq 1/\alpha}\alpha]\leq 0$.
Noting that $1-\inf_{\alpha:E^{\rm test}\geq 1/\alpha}\alpha=\sup_{\alpha:E^{\rm test}\geq 1/\alpha}\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}-\alpha=\sup_{\alpha>0}\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}-\alpha$, we obtain $\mathbb{E}[\sup_{\alpha>0}\mathbb{1}\{E^{\rm test}\geq 1/\alpha\}-\alpha]\leq 0$, i.e., $\sup_{\tilde{\alpha}>0}\mathbb{E}[\mathbb{1}\{E^{\rm test}\geq 1/\tilde{\alpha}\}-\tilde{\alpha}]\leq 0$.
Therefore, for any possibly data-dependent $\tilde{\alpha}>0$, $\mathbb{E}[\mathbb{1}\{E^{\rm test}\geq 1/\tilde{\alpha}\}]\leq\mathbb{E}[\tilde{\alpha}$], i.e., $\mathbb{P}(E^{\rm test}\geq 1/\tilde{\alpha})\leq\mathbb{E}[\tilde{\alpha}$]. The result then follows by substituting $E^{\rm test}$ and $\hat{C}_{n}^{\tilde{\alpha}}$ with their definitions. ∎